OpenStudy (anonymous):

What is the length of stack E F with bar on top when FG = 10 cm, BC = 8 cm, and AB = 9 cm? Round to the nearest tenth of a centimeter.

OpenStudy (anonymous):

Figure?

OpenStudy (anonymous):

one sec my computer died

OpenStudy (anonymous):

@eta

OpenStudy (anonymous):

Wanna start it from somewhere?

OpenStudy (anonymous):

Do we have options?

OpenStudy (anonymous):

can you start

OpenStudy (anonymous):

how do i start by setting it up @eta

OpenStudy (anonymous):

Find AC length first..

OpenStudy (anonymous):

I am not getting it actually but then also let us start..

OpenStudy (anonymous):

to find ac is is a^2+b^2=c^2

OpenStudy (anonymous):

I need to refresh on Properties of Right Pyramid..

OpenStudy (anonymous):

we can do a different one right now if you want

OpenStudy (anonymous):

Do we have options or answer choices?

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

for the other one yes

OpenStudy (anonymous):

@Nnesha

OpenStudy (anonymous):

Find AC, let us do it..

OpenStudy (anonymous):

\[AC^2 = AB^2 + BC^2\]

OpenStudy (anonymous):

12.04

OpenStudy (anonymous):

No. we are doing wrong I think.

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

A cone has a diameter of 30 in. and a slant height of 17 in. What is the approximate volume of the cone? Use 3.14 to approximate the value of straight pi. Round your answer to the nearest tenth, if necessary. (Hint: V equals 1 third straight pi r squared h ) A. 7536.0 in3 B. 5345.9 in3 C. 2419.9 in3 D. 1884.0 in3

OpenStudy (anonymous):

\[Volume = \frac{1}{3} \pi r^2 \cdot h\]

OpenStudy (anonymous):

r is radius here, but you are given SN = 30, SN is diameter of that circle, can you find Radius first?

OpenStudy (anonymous):

15

OpenStudy (anonymous):

it says diameter is 30 so radius is 15

OpenStudy (anonymous):

Good..

OpenStudy (anonymous):